the sum of three odd consecutive numbers equls 69
Found 2 solutions by user_dude2008, unlockmath:
Answer by user_dude2008(1862)

x+(x+2)+(x+4)=69
3x+6=69
3x=69-6
3x=63
x=21 <-------- First Number
x+2=23 <-------- second Number
x+4=25 <-------- 3rd Number

Answer by unlockmath(1688)

Hello,
Let's do this:
Have x be the first odd number.
Have x + 2 be the next odd number.
Have x + 4 be the next odd number.
We know all the above equals 69, right. So we can put it in the form of an equation.
x + (x + 2) + (x + 4) = 69
Then combine like terms.
3x + 6 = 69
Subtract 6 from both sides.
3x = 63
Divide both sides by 3.
x = 21
So now we know 21 is the first odd number and I'll let you figure out the next two odd numbers. Check the work to see if they all add up to 69.
RJ Toftness
Author "Unlock the Mystery to Math"
www.math-unlock.com